# HAFixAgent: History-Aware Automated Program Repair Agent

### v1

@@ 4.1.3. Historical Context
Once the blame commits are determined by the History Extractor, step 3 then mines relevant historical context from the identified blame commits by leveraging one of the three top-performing historical heuristics from our HAFix approach (Shi et al., 2025) (which focused on SL bugs only):
@@ 4.2. Agent Execution Loop
HAFixAgent’s execution loop follows an observe-act-feedback paradigm, inspired by frameworks like ReAct (Yao et al., 2023) and those used in recent software engineering agents (Yang et al., 2024). The loop proceeds with the following steps: 5 the LLM receives the current state (including the full conversation history of prior actions and observations); 6 the agent parses the model output to extract a tool invocation (i.e., a bash command, which the system prompt constrains the model always to provide) as the next action; 7 the command runs in a containerized tool sandbox that exposes repair utilities and dataset tools (e.g., search, edit and compile); 8 stdout and stderr are collected from the output of command execution as feedback; 9 compilation and test outcomes are captured and parsed; 10 the loop terminates on success or when a termination guard (e.g., step limit, cost limit, or timeout) is reached; otherwise, $1$ the observation is appended to the conversation history and fed back to the LLM for the next reasoning step.
@@ 5.1. Prompt Context Preparation
To prepare the non-history context as defined in Section 4.1.1, we need bug metadata such as bug reports, failing tests, and fault locations. We mine bug reports from issue tracker links provided by Defects4J. For 18 cases from the Chart project without a clear bug issue link, we use the fixing commit message as the bug description, and to prevent potential data leakage, we manually verified these commit messages to ensure they do not contain explicit fix instructions. Similar to the study (Nashid et al., 2025), we collect the failing tests provided by Defects4J. Consistent with common practices in the field of program repair (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025), we assume perfect fault localization and provide fault localization context, which is derived from the developer’s bug fix (following the process in Section 3.2), directly into the agent’s context. Specifically, similar to prior work (Bouzenia et al., 2025), this context consists of all file(s) and line(s) that need to be edited to fix the bug.
@@ 5.2. Baselines
We compare HAFixAgent with two recent APR techniques as baselines, i.e., RepairAgent (Bouzenia et al., 2025) and BIRCH-feedback (Nashid et al., 2025).
@@ 5.3. Metrics
In line with previous studies (Chen et al., 2021b; Parasaram et al., 2025; Ehsani et al., 2025; Nashid et al., 2025; Silva and Monperrus, 2025), we report Plausible@1 to evaluate the correctness based on whether the generated patch passes all test cases. Formally:
@@
Complementing Plausible@1, which measures the pass rate as a proportion of all bugs), we also report #Pass (the absolute count of bugs passing the test suite), similar to prior work (Nashid et al., 2025). To quantify the unique contribution of diverse historical context, we additionally report #Unique Pass, to show the number of bugs solved by a configuration using historical context but not by its non-history counterpart. We visualize overlaps using Venn diagrams to illustrate multi-way complementarity.
For efficiency, we report the average number of agent steps per bug, the average cost per bug in USD computed from token usage, and the cost stratified by outcome (successful versus failed fixes). To assess the statistical significance of differences in cost and number of steps, we compare the distributions across the four context configurations. Since the data is paired (same bugs, different configurations) and not assumed to be normally distributed, we employ the Friedman test (Friedman, 1937), a non-parametric test for multiple comparisons. If the Friedman test is significant ($p<0.05$), we conduct pairwise Wilcoxon signed-rank tests (Wilcoxon, 1992) for post-hoc analysis to compare each history-aware configuration against the non-history configuration. To control for multiple comparisons (three pairs), we apply the Bonferroni correction (Dunn, 1961) and set the significance threshold $\alpha$.
@@ 5.4. Implementation
HAFixAgent is implemented in Python and builds on mini-swe-agent (Yang et al., 2024) for the execution loop, which aligns with our lightweight design. While mini-swe-agent (and swe-agent) is designed naturally to support solving PR issue benchmarks such as SWE-Bench (Jimenez et al., 2024), it lacks the capability of historical context management (e.g., history localizing and retrieval), the repairing strategies and tools specifically for resolving Defects4J bugs. It is an open-source, lightweight framework with about 100 lines of core loop code, enabling seamless integration of historical context extraction and bash tools orchestration during runtime. We use DeepSeek-V3.2-Exp (DeepSeek-AI et al., 2024; DeepSeek-AI, 2025) via the official API with temperature set to 0.0, based on the official DeepSeek recommendation (DeepSeek, 2025). Prompts are rendered at runtime with jinja2.
@@ 6.1.1. Motivation
Our prior work (Shi et al., 2025) shows that injecting historical context improves LLM repair for single-line bugs under simple prompting. History carries heuristic signals that are hard to infer from a static snapshot: the fault introducing change, coevolving functions and files, and often intent hints in commit diffs and messages. RQ0’s finding that most Defects4J bugs, even multi-hunk ones, have only one single historical blame commit, suggests that historical context could also scale to APR techniques for more complex bug types.
@@ 6.2.1. Motivation
While RQ1 demonstrates that incorporating historical information into HAFixAgent enhances repair performance, it is also important to understand the associated cost and efficiency trade-offs. Cost is a first-order concern for agent systems because iterative tool use can accumulate substantial token and time budgets, even on simple bugs (Bouzenia et al., 2025; Xia and Zhang, 2024). Recent evaluations also highlight the risk of “token snowballs” and expensive failures (a phenomenon where unresolved attempts consume several times more tokens than successful ones (Fan et al., 2025), making cost and efficiency as important to measure as accuracy. HAFixAgent integrates additional contextual information into each prompt, which might potentially increase inference cost and agent reasoning steps. Therefore, we analyze whether the performance gains observed in RQ2 come at a significant cost and how efficiently the agent converges toward successful repairs across bug categories.
@@ 7.1. Performance Regressions
We observe 32 cases where a history configuration underperforms the non-history one or misses non-history fixes (Figure 3). For instance, in MFMH, the non-history configuration attains 50 while fn_all, fn_pair, and fl_diff reach 47, 43, and 48, respectively, even though each contributes 12 to 13 unique fixes that the non-history does not solve. This pattern aligns with LLM evidence that more context can hurt: models are easily distracted by irrelevant additions (Shi et al., 2023), and performance often declines when relevant facts sit in the middle or deeper parts of long prompts (Levy et al., 2024). In our setting, extra historical snippets can dilute salient context and reduce Plausible@1 on some bugs that don’t need historical context. Future work might filter and summarize history for direct relevance or add a guardrail that falls back to a leaner prompt when early steps show no progress.
@@ 7.3. Future Work
A promising, though more complex, frontier is to evolve HAFixAgent from a single-agent system into a collaborative multi-agent architecture (Rafi et al., 2024; Lin et al., 2025; Lee et al., 2024). Instead of one agent handling all tasks, we could design a system with specialized roles. For instance, a Coordinator agent could oversee the process, dispatching a specialized History agent to analyze repository patterns, blame commits, and issue reports to form a repair hypothesis. This hypothesis would then be passed to a Fixer agent focused on patch generation. Critically, a Reviewer agent could then critique the proposed patch, checking it not only against the test suite but also against the historical context for style, conventions, and potential side effects, which is similar to human code review. This separation of concerns could be particularly effective for the complex multi-file-multi-hunk (MFMH) bugs, where a single agent’s reasoning chain and context window can be overwhelmed.
@@ 8.1. Internal Validity
Our evaluation assumes perfect fault localization to isolate the evaluation of historical context on patch generation, a common setup in LLM-based APR studies (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025). Note that even given perfect fault localization, as most of the current work did, the LLM-based APR still struggle with complicated bugs (e.g., 39.1% Plausible for non-history configuration in MFMH bugs in our study). In practice, FL is noisy, especially for human-reported issues and multi-hunk defects, which can lower repair rates and lead to potentially wrong repair directions. For iterative agentic APR, updating FL during repair might help, but it does not eliminate the challenge. So we leverage perfect fault localization to isolate the evaluation of the historical context in HAFixAgent.
@@ 9.1. Traditional Automated Program Repair
Traditional APR formulates patch generation as a search or constraint problem over program transformations guided by tests or specifications. Search-based systems evolve patches with genetic or heuristic search over mutation operators and validate candidates against the test suite, exemplified by GenProg and successors (Le Goues et al., 2012; Le et al., 2016; Wen et al., 2018). Semantics or constraint-based repair derives patches by solving synthesis constraints obtained from symbolic execution or program analysis, as in SemFix (Nguyen et al., 2013) and Angelix (Mechtaev et al., 2016), and with targeted condition synthesis such as Nopol (Xuan et al., 2017). Template-based repair uses human-designed or mined fix schema that match buggy contexts and instantiate edits such as TBar (Liu et al., 2019b) and AVATAR (Liu et al., 2019a). A complementary line reuses in-project code fragments as ingredients, such as ssFix (Xin and Reiss, 2017) and CAPGEN (Wen et al., 2018). These approaches established core pipelines for localization, candidate generation, and test-based validation, but their coverage is constrained by operator or pattern design and by search scalability.
Learning-based APR before LLMs frames repair as code-to-code translation or edit prediction trained on a large amount of code corpus. Early neural systems include DeepFix (Gupta et al., 2017) for syntax errors and sequence-to-sequence repair, such as SequenceR (Chen et al., 2021a), with later improvements from model ensembling and syntax guidance in CoCoNuT (Lutellier et al., 2020), CURE (Jiang et al., 2021), and Recoder (Zhu et al., 2021). Template mining and ranking were also brought to production, for example, Facebook’s Getafix (Bader et al., 2019) for static-analyzer warnings and the end-to-end SapFix (Marginean et al., 2019) pipeline that proposed developer-reviewed patches. Together, traditional APR demonstrates effective repair across bug classes while revealing recurring limitations in computation cost, low accuracy and overfitting to tests. These limitations have motivated LLM and agent-based APR that learn fixes from data and plan tool-driven repair loops, which we introduce next.
@@ 9.2. LLM and Agent based Automated Program Repair
Large language models (LLMs) have significantly advanced Automated Program Repair (APR), moving from direct prompting to sophisticated, multi-step agentic systems. Early evidence shows code LLMs already outperform prior learning-based APR and benefit from fine-tuning and careful evaluation on standard benchmarks (Jiang et al., 2023; Xia et al., 2023b). This led to a variety of prompting and interaction strategies. For example, conversation-driven repair simulates a developer’s debugging process by interleaving patch generation with test feedback, achieving a strong balance of cost and accuracy (Xia and Zhang, 2024). Other approaches focus on improving the model’s reasoning, such as self-directed repair, which uses a chain-of-thought process to gather knowledge before fixing (Yin et al., 2024). The plastic surgery line revisits ingredient reuse by aligning LLMs with project-specific code (Xia et al., 2023a). Beyond pure prompting, hybrid and template-guided methods constrain or scaffold the patch space with analysis or repair templates to improve plausibility and correctness (Li et al., 2025a; Huang et al., 2025).
Agent-based APR systems represent a further step, shifting from single-shot patch generation to autonomous, tool-using workflows. Pioneering general-purpose software agents, like SWE-agent (Yang et al., 2024) and OpenHands (Wang et al., 2025), established a powerful paradigm by equipping LLMs with tools to interact with a repository, such as editors, shell commands, and test runners. Building on this, repair-focused agents have integrated more specialized tools. For instance, some employ search and fault localization to narrow down the buggy code (Bouzenia et al., 2025; Zhang et al., 2024c), while others explore multi-agent collaboration to divide the debugging task (Lee et al., 2024), with empirical studies beginning to map this rapidly evolving design space (Meng et al., 2024). A recent trend also focuses on improving agent performance on complex, repository-level tasks by incorporating memory and experience, allowing agents to learn from prior repair trajectories (Mu et al., 2025; Chen et al., 2025). These lines generally rely on local code and runtime outputs, with limited study of commit history as a first-class context inside the loop, which is the focus of our work.
@@ 9.3. In-context Learning for Automated Program Repair
A core question for LLM-based APR is which context to provide and how to structure it. Prior work shows that adding bug-related facts such as error messages, stack traces, and failing tests improves repair, and that the choice of facts matters (Parasaram et al., 2025; Xia and Zhang, 2024). Studies on local context indicate that models are sensitive to how much and which code surrounds the edit region (Prenner and Robbes, 2024). Other work leverages natural language bug reports or transforms descriptions into edits, and explores execution traces to guide patching (Fakhoury et al., 2024; Haque et al., 2025). Repository-level evaluations report that larger, realistic tasks remain challenging and require broader context curation (Chen et al., 2024).
Beyond local signals, historical context is receiving attention. HAFix (Shi et al., 2025) demonstrates that blame-derived commit data is an effective, lightweight historical signal for fixing more bugs, motivating history-aware agent designs at Defects4J scale. Layered knowledge injection systematically adds project and repository knowledge to prompts and improves bug fixing on broader-context scenarios (Ehsani et al., 2025). Built on top of HAFix, HAFixAgent advances this direction by bringing blame-driven history into the agentic workflow and quantifying its effect on single-hunk and multi-hunk bugs.

### v2

@@ 4.1.3. Historical Context
Once the blame commits are determined by the History Extractor, step 3 then mines relevant historical context from the identified blame commits by leveraging one of the three top-performing historical heuristics from our HAFix approach (shi2025hafix) (which focused on SL bugs only):
@@ 4.2. Agent Execution Loop
HAFixAgent’s execution loop follows an observe-act-feedback paradigm, inspired by frameworks like ReAct (yao2023react) and those used in recent software engineering agents (yang2024sweagent). The loop proceeds with the following steps: 5 the LLM receives the current state (including the full conversation history of prior actions and observations); 6 the agent parses the model output to extract a tool invocation (i.e., a bash command, which the system prompt constrains the model always to provide) as the next action; 7 the command runs in a containerized tool sandbox that exposes repair utilities and dataset tools (e.g., search, edit and compile); 8 stdout and stderr are collected from the output of command execution as feedback; 9 compilation and test outcomes are captured and parsed; 10 the loop terminates on success or when a termination guard (e.g., step limit, cost limit, or timeout) is reached; otherwise, $1$ the observation is appended to the conversation history and fed back to the LLM for the next reasoning step.
@@ 5.1. Prompt Context Preparation
To prepare the non-history context as defined in Section 4.1.1, we need bug metadata such as bug reports, failing tests, and fault locations. We mine bug reports from issue tracker links provided by Defects4J. For 18 cases from the Chart project without a clear bug issue link, we use the fixing commit message as the bug description, and to prevent potential data leakage, we manually verified these commit messages to ensure they do not contain explicit fix instructions. Similar to the study (nashid2025characterizing), we collect the failing tests provided by Defects4J. Consistent with common practices in the field of program repair (yang2024sweagent; fan2023automated; bouzenia2024repairagent; xia2022less; ye2024iter; nashid2025characterizing; ehsani2025bug), we assume perfect fault localization and provide fault localization context, which is derived from the developer’s bug fix (following the process in Section 3.2), directly into the agent’s context. Specifically, similar to prior work (bouzenia2024repairagent), this context consists of all file(s) and line(s) that need to be edited to fix the bug.
@@ 5.2. Baselines
We compare HAFixAgent with two recent APR techniques as baselines, i.e., RepairAgent (bouzenia2024repairagent) and BIRCH-feedback (nashid2025characterizing).
@@ 5.3. Metrics
In line with previous studies (chen2021evaluating; parasaram2024fact; ehsani2025bug; nashid2025characterizing; silva2025repairbench), we report Plausible@1 to evaluate the correctness based on whether the generated patch passes all test cases. Formally:
@@
Complementing Plausible@1, which measures the pass rate as a proportion of all bugs), we also report #Pass (the absolute count of bugs passing the test suite), similar to prior work (nashid2025characterizing). To quantify the unique contribution of diverse historical context, we additionally report #Unique Pass, to show the number of bugs solved by a configuration using historical context but not by its non-history counterpart. We visualize overlaps using Venn diagrams to illustrate multi-way complementarity.
For efficiency, we report the average number of agent steps per bug, the average cost per bug in USD computed from token usage, and the cost stratified by outcome (successful versus failed fixes). To assess the statistical significance of differences in cost and number of steps, we compare the distributions across the four context configurations. Since the data is paired (same bugs, different configurations) and not assumed to be normally distributed, we employ the Friedman test (friedman1937use), a non-parametric test for multiple comparisons. If the Friedman test is significant ($p<0.05$), we conduct pairwise Wilcoxon signed-rank tests (wilcoxon1992individual) for post-hoc analysis to compare each history-aware configuration against the non-history configuration. To control for multiple comparisons (three pairs), we apply the Bonferroni correction (dunn1961multiple) and set the significance threshold $\alpha$.
@@ 5.4. Implementation
HAFixAgent is implemented in Python and builds on mini-swe-agent (yang2024sweagent) for the execution loop, which aligns with our lightweight design. While mini-swe-agent (and swe-agent) is designed naturally to support solving PR issue benchmarks such as SWE-Bench (jimenez2023swebench), it lacks the capability of historical context management (e.g., history localizing and retrieval), the repairing strategies and tools specifically for resolving Defects4J bugs. It is an open-source, lightweight framework with about 100 lines of core loop code, enabling seamless integration of historical context extraction and bash tools orchestration during runtime. We use DeepSeek-V3.2-Exp (liu2024deepseek; deepseek-v3.2-exp-tech-note) via the official API with temperature set to 0.0, based on the official DeepSeek recommendation (deepseek-parameter-settings). Prompts are rendered at runtime with jinja2.
@@ 6.1.1. Motivation
Our prior work (shi2025hafix) shows that injecting historical context improves LLM repair for single-line bugs under simple prompting. History carries heuristic signals that are hard to infer from a static snapshot: the fault introducing change, coevolving functions and files, and often intent hints in commit diffs and messages. RQ0’s finding that most Defects4J bugs, even multi-hunk ones, have only one single historical blame commit, suggests that historical context could also scale to APR techniques for more complex bug types.
@@ 6.2.1. Motivation
While RQ1 demonstrates that incorporating historical information into HAFixAgent enhances repair performance, it is also important to understand the associated cost and efficiency trade-offs. Cost is a first-order concern for agent systems because iterative tool use can accumulate substantial token and time budgets, even on simple bugs (bouzenia2024repairagent; xia2024automated). Recent evaluations also highlight the risk of “token snowballs” and expensive failures (a phenomenon where unresolved attempts consume several times more tokens than successful ones (DBLP:journals/corr/abs-2509-09853), making cost and efficiency as important to measure as accuracy. HAFixAgent integrates additional contextual information into each prompt, which might potentially increase inference cost and agent reasoning steps. Therefore, we analyze whether the performance gains observed in RQ2 come at a significant cost and how efficiently the agent converges toward successful repairs across bug categories.
@@ 7.1. Performance Regressions
We observe 32 cases where a history configuration underperforms the non-history one or misses non-history fixes (Figure 3). For instance, in MFMH, the non-history configuration attains 50 while fn_all, fn_pair, and fl_diff reach 47, 43, and 48, respectively, even though each contributes 12 to 13 unique fixes that the non-history does not solve. This pattern aligns with LLM evidence that more context can hurt: models are easily distracted by irrelevant additions (shi2023large), and performance often declines when relevant facts sit in the middle or deeper parts of long prompts (levy2024same). In our setting, extra historical snippets can dilute salient context and reduce Plausible@1 on some bugs that don’t need historical context. Future work might filter and summarize history for direct relevance or add a guardrail that falls back to a leaner prompt when early steps show no progress.
@@ 7.3. Future Work
A promising, though more complex, frontier is to evolve HAFixAgent from a single-agent system into a collaborative multi-agent architecture (rafi2024multi; DBLP:conf/icse/LinKC25; lee2024unified). Instead of one agent handling all tasks, we could design a system with specialized roles. For instance, a Coordinator agent could oversee the process, dispatching a specialized History agent to analyze repository patterns, blame commits, and issue reports to form a repair hypothesis. This hypothesis would then be passed to a Fixer agent focused on patch generation. Critically, a Reviewer agent could then critique the proposed patch, checking it not only against the test suite but also against the historical context for style, conventions, and potential side effects, which is similar to human code review. This separation of concerns could be particularly effective for the complex multi-file-multi-hunk (MFMH) bugs, where a single agent’s reasoning chain and context window can be overwhelmed.
@@ 8.1. Internal Validity
Our evaluation assumes perfect fault localization to isolate the evaluation of historical context on patch generation, a common setup in LLM-based APR studies (yang2024sweagent; fan2023automated; bouzenia2024repairagent; xia2022less; ye2024iter; nashid2025characterizing; ehsani2025bug). Note that even given perfect fault localization, as most of the current work did, the LLM-based APR still struggle with complicated bugs (e.g., 39.1% Plausible for non-history configuration in MFMH bugs in our study). In practice, FL is noisy, especially for human-reported issues and multi-hunk defects, which can lower repair rates and lead to potentially wrong repair directions. For iterative agentic APR, updating FL during repair might help, but it does not eliminate the challenge. So we leverage perfect fault localization to isolate the evaluation of the historical context in HAFixAgent.
@@ 9.1. Traditional Automated Program Repair
Traditional APR formulates patch generation as a search or constraint problem over program transformations guided by tests or specifications. Search-based systems evolve patches with genetic or heuristic search over mutation operators and validate candidates against the test suite, exemplified by GenProg and successors (GenProg; le2016history; wen2018context). Semantics or constraint-based repair derives patches by solving synthesis constraints obtained from symbolic execution or program analysis, as in SemFix (nguyen2013semfix) and Angelix (mechtaev2016angelix), and with targeted condition synthesis such as Nopol (xuan2016nopol). Template-based repair uses human-designed or mined fix schema that match buggy contexts and instantiate edits such as TBar (liu2019tbar) and AVATAR (liu2019avatar). A complementary line reuses in-project code fragments as ingredients, such as ssFix (DBLP:conf/kbse/XinR17) and CAPGEN (wen2018context). These approaches established core pipelines for localization, candidate generation, and test-based validation, but their coverage is constrained by operator or pattern design and by search scalability.
Learning-based APR before LLMs frames repair as code-to-code translation or edit prediction trained on a large amount of code corpus. Early neural systems include DeepFix (gupta2017deepfix) for syntax errors and sequence-to-sequence repair, such as SequenceR (chen2019sequencer), with later improvements from model ensembling and syntax guidance in CoCoNuT (lutellier2020coconut), CURE (jiang2021cure), and Recoder (zhu2021syntax). Template mining and ranking were also brought to production, for example, Facebook’s Getafix (bader2019getafix) for static-analyzer warnings and the end-to-end SapFix (marginean2019sapfix) pipeline that proposed developer-reviewed patches. Together, traditional APR demonstrates effective repair across bug classes while revealing recurring limitations in computation cost, low accuracy and overfitting to tests. These limitations have motivated LLM and agent-based APR that learn fixes from data and plan tool-driven repair loops, which we introduce next.
@@ 9.2. LLM and Agent based Automated Program Repair
Large language models (LLMs) have significantly advanced Automated Program Repair (APR), moving from direct prompting to sophisticated, multi-step agentic systems. Early evidence shows code LLMs already outperform prior learning-based APR and benefit from fine-tuning and careful evaluation on standard benchmarks (jiang2023impact; xia2023automated). This led to a variety of prompting and interaction strategies. For example, conversation-driven repair simulates a developer’s debugging process by interleaving patch generation with test feedback, achieving a strong balance of cost and accuracy (xia2024automated). Other approaches focus on improving the model’s reasoning, such as self-directed repair, which uses a chain-of-thought process to gather knowledge before fixing (yin2024thinkrepair). The plastic surgery line revisits ingredient reuse by aligning LLMs with project-specific code (xia2023plastic). Beyond pure prompting, hybrid and template-guided methods constrain or scaffold the patch space with analysis or repair templates to improve plausibility and correctness (li2024hybrid; huang2025template).
Agent-based APR systems represent a further step, shifting from single-shot patch generation to autonomous, tool-using workflows. Pioneering general-purpose software agents, like SWE-agent (yang2024sweagent) and OpenHands (wang2024openhands), established a powerful paradigm by equipping LLMs with tools to interact with a repository, such as editors, shell commands, and test runners. Building on this, repair-focused agents have integrated more specialized tools. For instance, some employ search and fault localization to narrow down the buggy code (bouzenia2024repairagent; zhang2024autocoderover), while others explore multi-agent collaboration to divide the debugging task (lee2024unified), with empirical studies beginning to map this rapidly evolving design space (meng2024empirical). A recent trend also focuses on improving agent performance on complex, repository-level tasks by incorporating memory and experience, allowing agents to learn from prior repair trajectories (mu2025experepair; chen2025swe). These lines generally rely on local code and runtime outputs, with limited study of commit history as a first-class context inside the loop, which is the focus of our work.
@@ 9.3. In-context Learning for Automated Program Repair
A core question for LLM-based APR is which context to provide and how to structure it. Prior work shows that adding bug-related facts such as error messages, stack traces, and failing tests improves repair, and that the choice of facts matters (parasaram2024fact; xia2024automated). Studies on local context indicate that models are sensitive to how much and which code surrounds the edit region (prenner2024out). Other work leverages natural language bug reports or transforms descriptions into edits, and explores execution traces to guide patching (fakhoury2024nl2fix; haque2025towards). Repository-level evaluations report that larger, realistic tasks remain challenging and require broader context curation (chen2024large).
Beyond local signals, historical context is receiving attention. HAFix (shi2025hafix) demonstrates that blame-derived commit data is an effective, lightweight historical signal for fixing more bugs, motivating history-aware agent designs at Defects4J scale. Layered knowledge injection systematically adds project and repository knowledge to prompts and improves bug fixing on broader-context scenarios (ehsani2025bug). Built on top of HAFix, HAFixAgent advances this direction by bringing blame-driven history into the agentic workflow and quantifying its effect on single-hunk and multi-hunk bugs.